# Mastering AI: Big Data, Deep Learning, and the Evolution of Large Language Models - AutoML from Basics to State-of-the-Art Techniques

### v2

@@
In recent years, Artificial Intelligence (AI) and Machine Learning (ML) have grown tremendously in popularity across various industries. From healthcare and finance to retail and automotive, adopting machine learning models has led to significant advancements [1]. However, building machine learning models traditionally requires deep knowledge in multiple areas, such as data preprocessing, feature engineering, model selection, hyperparameter tuning, and evaluation [2]. For many beginners and even experienced practitioners, this process can be time-consuming and technically challenging.
This is where AutoML (Automated Machine Learning) comes in. AutoML simplifies the process of building machine learning models by automating many of the steps that would otherwise require manual intervention [3]. AutoML tools can automatically preprocess data, select the most suitable algorithms, and fine-tune hyperparameters to produce highly accurate models [4]. This automation not only speeds up the model development cycle but also allows users without deep knowledge of machine learning to create models with comparable performance to those made by experienced data scientists.
@@ 1.1.2 Boosting Productivity and Efficiency
With the automation of these processes, AutoML allows data scientists and machine learning engineers to focus on more critical tasks, such as understanding the business problem, interpreting the model results, and ensuring the ethical use of AI [3]. This increased productivity and efficiency can lead to faster deployment of models and, ultimately, more competitive advantages for organizations [5].
For example, in the healthcare industry, AutoML is being used to build models that can automatically diagnose medical conditions from data, such as detecting cancerous cells in X-ray images [6]. Such models, when deployed, can assist doctors in making faster, more accurate diagnoses [7].
Another example is in financial services. AutoML is being used by banks to develop models that detect fraudulent transactions in real-time, allowing financial institutions to save millions of dollars [8]. Such advances have already led to a demand for fewer manual fraud analysts, as machines are increasingly taking over these repetitive, pattern-based tasks [9].
@@ 1.1.3 Lowering the Entry Barrier for Beginners
AutoML makes machine learning more accessible to those without a strong background in the field. For beginners or newcomers to machine learning, it is now possible to build sophisticated models without needing to understand every intricate detail of the underlying algorithms [10]. This is especially beneficial for professionals from non-technical backgrounds who want to leverage machine learning in their work [11].
@@
18class SimpleNN(nn.Module):
@@
21 self.fc1 = nn.Linear(X_train.shape[1], 16)
22 self.fc2 = nn.Linear(16, 8)
23 self.fc3 = nn.Linear(8, 1)
@@ 2.1 Introduction to Python
Python [12, 13] is a high-level, interpreted programming language created by Guido van Rossum and first released in 1991 [14]. Its design philosophy emphasizes code readability, and its syntax allows programmers to express concepts in fewer lines of code than possible in languages like C++ or Java.
@@ 2.2 Install Python
In this section, we will guide you step by step through the installation of Python and setting up a suitable development environment. This includes installing IDLE [15], PyCharm [16], VSCode [17], and Anaconda [18], and setting up a virtual environment. These tools and environments are widely used for Python programming and offer distinct features beneficial for beginners.
@@ 2.9.2 Numpy
NumPy [19] is the fundamental package for scientific computing in Python. It provides support for arrays and matrices, along with a collection of mathematical functions to operate on these data structures. NumPy arrays are more efficient than Python lists, and they provide a more compact way of working with large amounts of data.
@@ 2.9.3 Pandas
Pandas [20] is a powerful library for data manipulation and analysis. It introduces two main data structures: Series and DataFrame. A Series is a one-dimensional array, while a DataFrame is a two-dimensional, table-like structure.
@@ 2.9.4 Matplotlib
Matplotlib [21] is a library used for data visualization in Python. It allows you to create a variety of static, animated, and interactive plots.
@@ 2.9.5 Scikit-learn
Scikit-learn [22] is one of the most popular libraries for building machine learning models. It provides efficient implementations of machine learning algorithms like linear regression, decision trees, clustering, and more.
@@ 2.9.6 PyTorch
PyTorch [23] is an open-source deep learning framework widely used for developing machine learning models, especially in the area of deep learning and neural networks. It is known for its flexibility and ease of use, and it provides automatic differentiation through its autograd feature.
@@
6class SimpleNN(nn.Module):
@@
9 self.fc1 = nn.Linear(1, 10)
10 self.fc2 = nn.Linear(10, 1)
@@ 2.9.7 TensorFlow
TensorFlow [24] is a well-known open-source deep learning framework developed by Google. It has been widely adopted for building and training machine learning models, particularly in production environments. TensorFlow provides both high-level and low-level APIs, offering flexibility and scalability for various tasks. In addition to its general applications, TensorFlow also supports the use of pre-trained models, making it a powerful tool for transfer learning and fine-tuning [25].
@@ 3.1.1 Supervised Learning
Supervised learning is the most common form of machine learning. In this type, the model is trained using a labeled dataset, meaning that each input comes with an associated output. The goal of the algorithm is to learn the relationship between inputs and outputs in such a way that it can predict the output for new, unseen data [26]. Supervised learning is typically used in applications like classification (e.g., spam detection in emails [27]) and regression (e.g., predicting housing prices) [28].
@@
7class SimpleNN(nn.Module):
@@
10 self.layer1 = nn.Linear(2, 1) # Input: 2 features, Output: 1 (binary class)
@@
21criterion = nn.BCELoss() # Binary Cross Entropy Loss for classification
@@ 3.1.2 Unsupervised Learning
In unsupervised learning, the model is trained on data that has no labels. The goal is to uncover hidden patterns or structures within the data [29]. A common application of unsupervised learning is clustering, where the model groups similar data points together [30]. Another example is dimensionality reduction, where the model reduces the number of features in the dataset while retaining essential information [31].
@@ 3.1.3 Reinforcement Learning
Reinforcement learning is different from both supervised and unsupervised learning. In reinforcement learning, an agent interacts with an environment and learns by receiving feedback in the form of rewards or penalties [32]. The agent takes actions to maximize cumulative rewards over time. Applications include game playing, robotics, and self-driving cars [33].
Example: Imagine a robot learning to navigate a maze. Each time the robot takes a step, it receives a reward if it moves closer to the goal and a penalty if it moves further away. The robot continues to explore and adjust its actions based on the rewards and penalties received, with the ultimate aim of finding the shortest path to the goal [34].
@@ 3.2.1 Supervised Learning: A Detailed Look
6class Classifier(nn.Module):
@@
9 self.layer1 = nn.Linear(1, 1)
@@ 3.3.1 Accuracy, Precision, Recall, and F1-score
Accuracy is the simplest performance metric. It is the ratio of correctly predicted instances to the total number of instances. However, accuracy may not always be a good metric, especially in cases where the classes are imbalanced [35].
Precision is the ratio of true positive predictions to the total number of positive predictions (both true and false positives). It answers the question: ”Of all the instances predicted as positive, how many were actually positive?” [36].
Recall (also called sensitivity) is the ratio of true positive predictions to the total number of actual positive instances. It answers the question: ”Of all the actual positive instances, how many did the model correctly predict?” [36].
F1-score is the harmonic mean of precision and recall. It provides a balanced measure that takes both false positives and false negatives into account. It is especially useful when the class distribution is imbalanced [37].
@@ 3.3.2 ROC Curve and AUC
The ROC curve (Receiver Operating Characteristic curve) is a graphical representation of the performance of a binary classifier across different threshold values [38]. It is particularly useful for understanding the trade-off between two important metrics: the true positive rate (recall or sensitivity) and the false positive rate (1-specificity). The ROC curve plots the true positive rate on the y-axis and the false positive rate on the x-axis.
The ROC curve has its origins in World War II, where it was first used by radar operators to detect enemy objects. The operators had to balance the detection of real objects (true positives) against false alarms (false positives). This led to the development of the ROC curve to evaluate how well different detection strategies worked under varying conditions [39]. Over time, this concept was adapted for evaluating binary classification models in machine learning and medical testing [40].
@@
  • An AUC of 0.5 indicates that the classifier performs no better than random guessing [41].
@@ 4.3.1 Feature Selection
  • Correlation-based selection: Select features that have high correlation with the target variable but low correlation with each other [42].
  • Recursive Feature Elimination (RFE): Iteratively remove less important features and evaluate model performance [43].
  • Tree-based methods: Use the importance scores generated by tree-based models like Random Forests or Gradient Boosting to select features [44, 45].
@@ 5.1 Basic Principles of Linear Regression
Linear regression is a fundamental machine learning algorithm used for predicting a continuous target variable based on one or more input variables (also called features) [46]. The core idea is to model the relationship between the input variables and the output variable as a linear combination of the input features.
@@ 5.2 Ordinary Least Squares
The most common method for estimating the coefficients $\mathbf{w}$ and $b$ in linear regression is called Ordinary Least Squares (OLS). The idea behind OLS is to minimize the sum of the squared differences between the predicted values $\mathbf{\hat{y}}=\mathbf{Xw}+b$ and the actual target values $\mathbf{y}$ [47].
@@ 5.3.1 Ridge Regression
Ridge regression adds a penalty term to the loss function, which helps constrain the size of the weights and thus reduces the model’s complexity [48]. The modified loss function for Ridge regression is:
@@ 5.3.2 Lasso Regression
Lasso regression is another regularization technique that adds a penalty based on the L1 norm of the weights. The loss function is modified as follows [49]:
@@ 5.4 Implementation of Linear Regression
11class LinearRegressionModel(nn.Module):
@@
14 self.linear = nn.Linear(1, 1) # 1 input, 1 output
@@ 6.1 Basic Concepts of SVM
Support Vector Machines (SVM) are one of the most powerful and widely-used supervised machine learning algorithms for classification and regression problems [50]. SVM aims to find the optimal hyperplane that separates the data into distinct classes. In simpler terms, the algorithm looks for the best boundary (or decision surface) between the classes. The core idea is to maximize the margin (the distance between the decision boundary and the closest data points, called support vectors) while correctly classifying the data [51].
An SVM constructs a hyperplane or a set of hyperplanes in a high-dimensional space [52]. The key principle is that the hyperplane that maximizes the margin between the data points of different classes is the best choice [53].
@@
[…]
@@ 6.2.1 Linear SVM
In cases where the data is linearly separable (i.e., a straight line can separate the classes), a Linear SVM is sufficient. Linear SVM works well for simple datasets where the relationship between the input features and the output labels is linear [54].
@@ 6.2.2 Non-linear SVM
When the data is not linearly separable, we need a more complex boundary. In such cases, Non-linear SVM can be used, which employs the ”kernel trick” to transform the data into a higher-dimensional space where it becomes linearly separable [53].
@@ 6.3.1 Linear Kernel
The Linear Kernel is the simplest kernel, equivalent to the dot product between two vectors. It is suitable for linearly separable data [55].
@@ 6.3.2 Polynomial Kernel
The Polynomial Kernel allows learning more complex decision boundaries by introducing polynomial features. This is useful when the data is not linearly separable [50].
@@ 6.3.3 Radial Basis Function (RBF) Kernel
The RBF Kernel is the most commonly used kernel in practice because it can handle both linear and non-linear data. It maps the data to an infinite-dimensional space [56].
@@ 6.4 Implementation of SVM
29 self.linear = nn.Linear(2, 1) # 2 input features, 1 output
@@ 7.1 Basic Principles of Decision Trees
Decision trees are a popular and powerful machine learning algorithm used for both classification and regression tasks [57]. They work by splitting data into subsets based on the feature values. Each decision in the tree represents a condition on one feature, and the process of splitting continues until the tree reaches a state where further splitting does not significantly improve the model [58]. Decision trees are intuitive, easy to understand, and suitable for both small and large datasets [59].
@@ 7.1.1 How Decision Trees Work
A decision tree consists of nodes, where each node represents a decision based on a feature of the data. The tree starts with a root node, and branches are created based on the values of features. Each branch leads to either another decision node or a leaf node, where a final classification or regression value is predicted [60]. Let us visualize a simple decision tree:
[…]

### v3

@@
   Xinyuan Song
Emory University
xsong30@emory.edu
@@
In recent years, Artificial Intelligence (AI) and Machine Learning (ML) have grown tremendously in popularity across various industries. From healthcare and finance to retail and automotive, adopting machine learning models has led to significant advancements [jordan2015machine]. However, building machine learning models traditionally requires deep knowledge in multiple areas, such as data preprocessing, feature engineering, model selection, hyperparameter tuning, and evaluation [kohavi1998feature]. For many beginners and even experienced practitioners, this process can be time-consuming and technically challenging.
This is where AutoML (Automated Machine Learning) comes in. AutoML simplifies the process of building machine learning models by automating many of the steps that would otherwise require manual intervention [he2021automl]. AutoML tools can automatically preprocess data, select the most suitable algorithms, and fine-tune hyperparameters to produce highly accurate models [feurer2015efficient]. This automation not only speeds up the model development cycle but also allows users without deep knowledge of machine learning to create models with comparable performance to those made by experienced data scientists.
@@ 1.1.2 Boosting Productivity and Efficiency
With the automation of these processes, AutoML allows data scientists and machine learning engineers to focus on more critical tasks, such as understanding the business problem, interpreting the model results, and ensuring the ethical use of AI [he2021automl]. This increased productivity and efficiency can lead to faster deployment of models and, ultimately, more competitive advantages for organizations [zoller2021benchmark].
For example, in the healthcare industry, AutoML is being used to build models that can automatically diagnose medical conditions from data, such as detecting cancerous cells in X-ray images [esteva2019guide]. Such models, when deployed, can assist doctors in making faster, more accurate diagnoses [topol2019high].
Another example is in financial services. AutoML is being used by banks to develop models that detect fraudulent transactions in real-time, allowing financial institutions to save millions of dollars [radovanovic2020fraud]. Such advances have already led to a demand for fewer manual fraud analysts, as machines are increasingly taking over these repetitive, pattern-based tasks [pawar2020fraud].
@@ 1.1.3 Lowering the Entry Barrier for Beginners
AutoML makes machine learning more accessible to those without a strong background in the field. For beginners or newcomers to machine learning, it is now possible to build sophisticated models without needing to understand every intricate detail of the underlying algorithms [yao2018taking]. This is especially beneficial for professionals from non-technical backgrounds who want to leverage machine learning in their work [gonzalez2020helping].
@@
18class SimpleNN($n$.Module):
@@
21 self.fc1 = $n$.Linear(X_train.shape[1], 16)
22 self.fc2 = $n$.Linear(16, 8)
23 self.fc3 = $n$.Linear(8, 1)
@@ 2.1 Introduction to Python
Python [peng2024deeplearningmachinelearning, li2024deeplearningmachinelearning] is a high-level, interpreted programming language created by Guido van Rossum and first released in 1991 [van1995python]. Its design philosophy emphasizes code readability, and its syntax allows programmers to express concepts in fewer lines of code than possible in languages like C++ or Java.
@@ 2.2 Install Python
In this section, we will guide you step by step through the installation of Python and setting up a suitable development environment. This includes installing IDLE [kaswan2023python], PyCharm [islam2015mastering], VSCode [del2019visual], and Anaconda [phelps2000anaconda], and setting up a virtual environment. These tools and environments are widely used for Python programming and offer distinct features beneficial for beginners.
@@ 2.9.2 Numpy
NumPy [numpy] is the fundamental package for scientific computing in Python. It provides support for arrays and matrices, along with a collection of mathematical functions to operate on these data structures. NumPy arrays are more efficient than Python lists, and they provide a more compact way of working with large amounts of data.
@@ 2.9.3 Pandas
Pandas [pandas] is a powerful library for data manipulation and analysis. It introduces two main data structures: Series and DataFrame. A Series is a one-dimensional array, while a DataFrame is a two-dimensional, table-like structure.
@@ 2.9.4 Matplotlib
Matplotlib [matplotlib_] is a library used for data visualization in Python. It allows you to create a variety of static, animated, and interactive plots.
@@ 2.9.5 Scikit-learn
Scikit-learn [pedregosa2011scikit] is one of the most popular libraries for building machine learning models. It provides efficient implementations of machine learning algorithms like linear regression, decision trees, clustering, and more.
@@ 2.9.6 PyTorch
PyTorch [pytorch_] is an open-source deep learning framework widely used for developing machine learning models, especially in the area of deep learning and neural networks. It is known for its flexibility and ease of use, and it provides automatic differentiation through its autograd feature.
@@
6class SimpleNN($n$.Module):
@@
9 self.fc1 = $n$.Linear(1, 10)
10 self.fc2 = $n$.Linear(10, 1)
@@ 2.9.7 TensorFlow
TensorFlow [tensorflow2015-whitepaper] is a well-known open-source deep learning framework developed by Google. It has been widely adopted for building and training machine learning models, particularly in production environments. TensorFlow provides both high-level and low-level APIs, offering flexibility and scalability for various tasks. In addition to its general applications, TensorFlow also supports the use of pre-trained models, making it a powerful tool for transfer learning and fine-tuning [chen2024deeplearningmachinelearning].
@@ 3.1.1 Supervised Learning
Supervised learning is the most common form of machine learning. In this type, the model is trained using a labeled dataset, meaning that each input comes with an associated output. The goal of the algorithm is to learn the relationship between inputs and outputs in such a way that it can predict the output for new, unseen data [friedman2001elements]. Supervised learning is typically used in applications like classification (e.g., spam detection in emails [kotsiantis2007supervised]) and regression (e.g., predicting housing prices) [hastie2009elements].
@@
7class SimpleNN($n$.Module):
@@
10 self.layer1 = $n$.Linear(2, 1) # Input: 2 features, Output: 1 (binary class)
@@
21criterion = $n$.BCELoss() # Binary Cross Entropy Loss for classification
@@ 3.1.2 Unsupervised Learning
In unsupervised learning, the model is trained on data that has no labels. The goal is to uncover hidden patterns or structures within the data [ghahramani2004unsupervised]. A common application of unsupervised learning is clustering, where the model groups similar data points together [jain1999data]. Another example is dimensionality reduction, where the model reduces the number of features in the dataset while retaining essential information [van2009dimensionality].
@@ 3.1.3 Reinforcement Learning
Reinforcement learning is different from both supervised and unsupervised learning. In reinforcement learning, an agent interacts with an environment and learns by receiving feedback in the form of rewards or penalties [sutton2018reinforcement]. The agent takes actions to maximize cumulative rewards over time. Applications include game playing, robotics, and self-driving cars [mnih2015human].
Example: Imagine a robot learning to navigate a maze. Each time the robot takes a step, it receives a reward if it moves closer to the goal and a penalty if it moves further away. The robot continues to explore and adjust its actions based on the rewards and penalties received, with the ultimate aim of finding the shortest path to the goal [kaelbling1996reinforcement].
@@ 3.2.1 Supervised Learning: A Detailed Look
6class Classifier($n$.Module):
@@
9 self.layer1 = $n$.Linear(1, 1)
@@ 3.3.1 Accuracy, Precision, Recall, and F1-score
Accuracy is the simplest performance metric. It is the ratio of correctly predicted instances to the total number of instances. However, accuracy may not always be a good metric, especially in cases where the classes are imbalanced [he2009learning].
Precision is the ratio of true positive predictions to the total number of positive predictions (both true and false positives). It answers the question: ”Of all the instances predicted as positive, how many were actually positive?” [sokolova2006beyond].
Recall (also called sensitivity) is the ratio of true positive predictions to the total number of actual positive instances. It answers the question: ”Of all the actual positive instances, how many did the model correctly predict?” [sokolova2006beyond].
F1-score is the harmonic mean of precision and recall. It provides a balanced measure that takes both false positives and false negatives into account. It is especially useful when the class distribution is imbalanced [goutte2005probabilistic].
@@ 3.3.2 ROC Curve and AUC
The ROC curve (Receiver Operating Characteristic curve) is a graphical representation of the performance of a binary classifier across different threshold values [fawcett2006introduction]. It is particularly useful for understanding the trade-off between two important metrics: the true positive rate (recall or sensitivity) and the false positive rate (1-specificity). The ROC curve plots the true positive rate on the y-axis and the false positive rate on the x-axis.
The ROC curve has its origins in World War II, where it was first used by radar operators to detect enemy objects. The operators had to balance the detection of real objects (true positives) against false alarms (false positives). This led to the development of the ROC curve to evaluate how well different detection strategies worked under varying conditions [swets1979receiver]. Over time, this concept was adapted for evaluating binary classification models in machine learning and medical testing [hanley1982meaning].
@@
  • An AUC of 0.5 indicates that the classifier performs no better than random guessing [bradley1997use].
@@ 4.3.1 Feature Selection
  • Correlation-based selection: Select features that have high correlation with the target variable but low correlation with each other [hall1999correlation].
  • Recursive Feature Elimination (RFE): Iteratively remove less important features and evaluate model performance [guyon2002gene].
  • Tree-based methods: Use the importance scores generated by tree-based models like Random Forests or Gradient Boosting to select features [breiman2001random, chen2016xgboost].
@@ 5.1 Basic Principles of Linear Regression
Linear regression is a fundamental machine learning algorithm used for predicting a continuous target variable based on one or more input variables (also called features) [montgomery2012introduction]. The core idea is to model the relationship between the input variables and the output variable as a linear combination of the input features.
@@ 5.2 Ordinary Least Squares
The most common method for estimating the coefficients $\mathbf{w}$ and $b$ in linear regression is called Ordinary Least Squares (OLS). The idea behind OLS is to minimize the sum of the squared differences between the predicted values $\mathbf{\hat{y}}=\mathbf{Xw}+b$ and the actual target values $\mathbf{y}$ [seber2003linear].
@@ 5.3.1 Ridge Regression
Ridge regression adds a penalty term to the loss function, which helps constrain the size of the weights and thus reduces the model’s complexity [hoerl1970ridge]. The modified loss function for Ridge regression is:
@@ 5.3.2 Lasso Regression
Lasso regression is another regularization technique that adds a penalty based on the L1 norm of the weights. The loss function is modified as follows [tibshirani1996regression]:
@@ 5.4 Implementation of Linear Regression
11class LinearRegressionModel($n$.Module):
@@
14 self.linear = $n$.Linear(1, 1) # 1 input, 1 output
@@ 6.1 Basic Concepts of SVM
Support Vector Machines (SVM) are one of the most powerful and widely-used supervised machine learning algorithms for classification and regression problems [cortes1995support]. SVM aims to find the optimal hyperplane that separates the data into distinct classes. In simpler terms, the algorithm looks for the best boundary (or decision surface) between the classes. The core idea is to maximize the margin (the distance between the decision boundary and the closest data points, called support vectors) while correctly classifying the data [burges1998tutorial].
An SVM constructs a hyperplane or a set of hyperplanes in a high-dimensional space [vapnik1998statistical]. The key principle is that the hyperplane that maximizes the margin between the data points of different classes is the best choice [scholkopf2002learning].
@@
$\text{Maximize Margin:}\quad\frac{2}{\|\mathbf{w}\|}$
@@ 6.2.1 Linear SVM
In cases where the data is linearly separable (i.e., a straight line can separate the classes), a Linear SVM is sufficient. Linear SVM works well for simple datasets where the relationship between the input features and the output labels is linear [vapnik1995nature].
@@ 6.2.2 Non-linear SVM
When the data is not linearly separable, we need a more complex boundary. In such cases, Non-linear SVM can be used, which employs the ”kernel trick” to transform the data into a higher-dimensional space where it becomes linearly separable [scholkopf2002learning].
@@ 6.3.1 Linear Kernel
The Linear Kernel is the simplest kernel, equivalent to the dot product between two vectors. It is suitable for linearly separable data [hearst1998support].
@@ 6.3.2 Polynomial Kernel
The Polynomial Kernel allows learning more complex decision boundaries by introducing polynomial features. This is useful when the data is not linearly separable [cortes1995support].
@@ 6.3.3 Radial Basis Function (RBF) Kernel
The RBF Kernel is the most commonly used kernel in practice because it can handle both linear and non-linear data. It maps the data to an infinite-dimensional space [scholkopf1997comparing].
@@ 6.4 Implementation of SVM
29 self.linear = $n$.Linear(2, 1) # 2 input features, 1 output
@@ 7.1 Basic Principles of Decision Trees
Decision trees are a popular and powerful machine learning algorithm used for both classification and regression tasks [quinlan1986induction]. They work by splitting data into subsets based on the feature values. Each decision in the tree represents a condition on one feature, and the process of splitting continues until the tree reaches a state where further splitting does not significantly improve the model [breiman1984classification]. Decision trees are intuitive, easy to understand, and suitable for both small and large datasets [loh2011classification].
@@ 7.1.1 How Decision Trees Work
A decision tree consists of nodes, where each node represents a decision based on a feature of the data. The tree starts with a root node, and branches are created based on the values of features. Each branch leads to either another decision node or a leaf node, where a final classification or regression value is predicted [safavian1991survey]. Let us visualize a simple decision tree:
Is Temperature ¿ 30°C?YesIs Humidity ¿ 70%?Yes, Play Tennis = NoNo, Play Tennis = YesNo, Play Tennis = Yes